OpenStudy (anonymous):

What is the centripetal force of an object undergoing uniform circular motion when its radius is doubled and its speed remains constant? A. the same as before B. twice as great as before C. half as great as before D. four times as great as before

OpenStudy (ivancsc1996):

Well the equation is\[F _{c}=\frac{ mv ^{2} }{ r } \rightarrow F _{c} \alpha r ^{-1} = F _{c}\alpha \frac{ 1 }{ r }\] Which is the same as saying that as the radius (r) gets bigger, the centripetal force (Fc) gets smaller by the same amount. So if the radius is dobled then the centripetal force is half of the original. So the answer is: C. half as great as before

OpenStudy (anonymous):

i love you :3 thanks! :D